OpenStudy (anonymous):

Determine the relationship between the lines that pass through the given points: Line A: (-3, 5) and (0, 7) Line B: (6, 2) and (9, 4) A. they are perpendicular B. they are parallel C. they are the same line D. they have no relations

OpenStudy (anonymous):

is this dot product or cross product ?

OpenStudy (anonymous):

(-3 * 0) - (7 * 5) = -35 (6 * 9) - ( 4 * 2) = 46 thats there cross product.

OpenStudy (anonymous):

i think they want a cross product

OpenStudy (anonymous):

perpendicular is the sum of the cross product is = 0